▾▴ Toggle Texts
▾▴ Toggle Images
◱ Full Size Pics
⊻ Bottom of Page
- (inserted: 24.06.2012 20:00)
The first Sampling Campaign of the Dead Sea ICDP project ...
The first Sampling Campaign of the Dead Sea ICDP project began this week. Scientists and students from Israel, Germany and the USA arrived, while many more are expected next week, including researchers from Japan and Switzerland. Close to 100 m of core have been sampled so far, for many different projects, such as climatic, paleoseismic, lacustrine, biological, and other studies.
- (inserted: 27.03.2011 20:00)
Yesterday the barge was towed back to Arugot (Gonen station) ...
Yesterday the barge was towed back to Arugot (Gonen station) and today the big crane was working on lifting the rig etc. We hope to complete this phase of de-mobilization in a couple of days.
- (inserted: 24.03.2011 22:00)
And now the ICDP Dead Sea project comes to an ...And now the ICDP Dead Sea project comes to an end. All in all, we have two long cores - one from the center of the lake (455.48 mbsf) and one from near the southwestern shore (346.07 mbsf), and many short cores of a few tens of meters. Jochem and his logging team have finished their work and we are waiting to tow the barge back to Sylvie Gonen's base camp for demobilization. It has been an exciting project that will shed light on some major scientific questions. From all of us here in Israel, thank you and see you in Bremen!
- (inserted: 19.03.2011 22:00)
Coring at Site 3 is going well. We are already ...Coring at Site 3 is going well. We are already at -290m below the lake floor after a few days at work. Some problems were encountered - clogged drill bit, salt or sand that led to a drop in recovery, but the team soon had things under control again and were pulling up core after core. The sun is shining and it is hot in Ein Gedi!
- (inserted: 28.02.2011 22:00)
And we're back! After suspending operations in order to fix ...And we're back! After suspending operations in order to fix the barge, we are set to resume drilling tomorrow (Wednesday). The new site is in very shallow water, which will of course, make things easier. We are expecting a slightly different lithology here as we are very close to the western shore. Perhaps a little more salt in the beginning. Perhaps lower rates of sedimentation. Looking forward to a successful second campaign!
- (inserted: 17.01.2011 22:00)
After a long break while we waited for parts to ...After a long break while we waited for parts to arrive from the States, we tried to resume operations over the weekend. However, it seems that the problems were not fixed, despite the efforts made by DOSECC and the drill team. A decision was made to temporarily shutdown. Coring at Site 2 will resume on March 1st and continue for around 2 weeks. During the interim period, the engineer that built the rig will be flown in to work out all the bugs. We would like to take this opportunity to thank Beau Marshall, Steve Cole, Rich Szentmiklosi, A.J. Vecchiarelli, Joe Bolin, Joe Hammond and Eli Driggers for their tremendous dedication, hard work and for helping us achieve our scientific goals. We hope to see you back for Stage 2. Special thanks also goes to Sylvie Gonen and the maritime logistic support team of the RV Taglit: Gideon Amit, Dubi Klein, Ofir Zukerman; as well as our spokesperson Avital Baer, the staff at the Ein Gedi Guest House and at the Kibbutz; Yael Dagoni, Elitza Hadzhiivanova, Elisa Kagan, Lisa Coianiz, Ina Neugebauer, Emily Goldstein, Jason Jweda, Paul Neitzel, the enthusiastic student volunteers; the PIs and scientists: Amotz Agnon, Daniel Aritzegui, Achim Brauer, Steve Goldstein, Uli Harms, Gerlad Haug, Emi Ito, Hiroyuki Kitagawa, Yoshinori Yasuda, Yossi Yechieli, Nicolas Waldmann; the logging team: Jochem Kueck, Matxalen Rey Abasolo, Martin Toepfer, and Christian Carnein; and other visitors and friends who put in time and helped us along the way. The list goes on and on ...
- (inserted: 06.01.2011 22:00)
We have suspended operations until Wednesday while we wait for ...We have suspended operations until Wednesday while we wait for a part to arrive from the States. The weather has turned very dramatic. The sky is purple-gray with rays of light shining through. Last night it poured. As I walk across the salt encrusted beach, where very few (if any) venture, I am amazed at what I find. From dead bugs to strange formations. I leave you for a few days with a picture of hope...
- (inserted: 04.01.2011 22:00)
The barge was towed yesterday to the vicinity of Site ...The barge was towed yesterday to the vicinity of Site 2 and anchored today. It was a long, hard process, but everyone pitched in to help. A very nice example of cooperation between the science crew and the drillers, with PIs Steve Goldstein and Moti Stein and Project Manager Michael Lazar jumping in the water to help Beau Marshall drag the anchor cable to the post onshore. There is a good feeling of expectation as we begin to core again with the night shift.
- (inserted: 02.01.2011 22:00)
Coring is going well. We are at about 375 m ...Coring is going well. We are at about 375 m below the lake floor (~75m into the sedimentary section) with the target of this hole being ~89 m - i.e. the end of the Holocene and the beginning of the Lisan formation. Recovery has been about 80%. I personally built 4 anchors today with the help of the machine-shop at Kibbutz Ein Gedi. The plan is now to pull up tomorrow and move to Site 2 - opposite the Ein Gedi spa, close to shore (200 m) in a water depth of 20 m. Drilling should go smoother and since it is in a small bay, the lake should be calmer. We will keep our fingers crossed.
- (inserted: 30.12.2010 22:00)
As of 16:00 today, we are at 354 m down ...As of 16:00 today, we are at 354 m down in Hole H. Last night they finished drilling without coring to about 40 meters beneath the lake floor to get past the problematic layer. Cores started to come out again, which is good news. The weather is overcast and cloudy. Rain has been falling on and off all day, and it is very dramatic (see picture of the day) but the lake has been calm. Wishing you all a Happy New Year and a great 2011.
- (inserted: 28.12.2010 22:00)
More problems. The weather has not been good to us. ...More problems. The weather has not been good to us. The winds and currents have pushed us off-hole and bent the drill pipe.In addition, salt accumulation on the anchor cables is making them heavy and causing problems. After tripping up again, moving to Hole G (!) and tripping down, we are ready to begin once more, but... it was too windy to change the shifts. Night crew were stuck out on the platform again. Frustration has set in and the mood around here is not good.
- (inserted: 26.12.2010 22:00)
Christmas came and went. We decided to give the drillers ...Christmas came and went. We decided to give the drillers a little break and shut down operations for 24 hours so they could go to Bethlehem. We hope that it boosted their morale as they were feeling a little worn out and disappointed at loosing the bottom hole assembly again. We are now in hole F. The plan is to drill down to 45 m below the lake floor to get past the problematic unstable zone in the 30's. They will try to extract core using the Extended Nose tool until the container is released from customs. There were some problems with retrieving the non-coring tool during one of the runs yesterday, but the patience and expertise of the drillers paid off and they pulled it up. Bad weather kept the night shift out on the platform again today. Hopefully, they slept well during the day and are ready to get back to work. We just sent out a boat to exchange the science crew.
- (inserted: 24.12.2010 22:00)
Once again, we lost the bit. Something is going on ...Once again, we lost the bit. Something is going on at around 33 m into the sediment. Beau explained that 6 pipes were in the hole when it broke meaning that it failed at the interface between the water and the lake floor. This could be caused by the salt falling around the pipe and gripping it, a deviated pipe from the rig to the seafloor (through the water column) caused by the rig drifting off-hole, or a combination of both. As this was the very last inner-tube for the alien and the HPC, drilling came to a halt. The container with the spare parts is in Israel, but being held up by bureaucratic issues at customs. We gave the drillers 24 hours off to go to Bethlehem for Christmas, which really improved their morale. They went back out today at 13:00 and are planning to trip down to Hole F and begin using the "extended bit" - which can only go through soft sediments (not salt). Hopefully, a Christmas miracle will occur and the container will be released and on its way tomorrow. Merry Christmas from all of us here!
- (inserted: 23.12.2010 22:00)
Another inner-tube was bent when shooting the HPC in ...Another inner-tube was bent when shooting the HPC in Hole D. The drillers think that it may be due to pressurizing the tool too much after it fires. Luckily, they had another emergency tube that was brought from Lake Van, which had a small piece of core trapped in it (very different from the sediments found here!). The weather did not permit a shift change yesterday morning, but a new science team was sent out for the night shift. We brought a small Christmas tree with us and set it up on the barge in front of the drillers' shack - a little token of our appreciation. Since we had to pull out of Hole D, we began Hole E, but 3 meters below the lake floor, in the sediment. HPC was fired without driving it from above with pressure and recovery was good. The day shift reports that they too are getting good runs. We are keeping our fingers crossed.
- (inserted: 20.12.2010 22:00)
There has been a little feeling of frustration in the ...There has been a little feeling of frustration in the camp the last few days. Basically since the storm we have not been able to get good recovery. Hole B was lost and Hole C was not going as expected. All indication led Beau and his team to the conclusion the there was still bent pipe in the hole - damage from the storm or from the pressure of waves during the last few days. The loggers went out yesterday during the day shift and confirmed that the deviation was more than 11 degrees (!!) in some sections. A decision was made to pull out, replace all 60 lengths of pipe and begin again in Hole D. This time, the plan is take is slowly and really try to get good recovery for the top 100 m.
- (inserted: 18.12.2010 22:00)
For some reason, coring is not going so well and ...For some reason, coring is not going so well and the drillers are constantly battling to recover anything from the third hole. Apparently, there was also some cave-in this time around. Each run conducted they try something new - changing the core-catcher, the drill bit, adjusting feed, rotation, etc. Nothing seems to be working. Beau went out with the morning shift to try to disassemble the sampling tool and check if there is a problem with it. It could also be that the formation her is much different than that found in the first hole and is the reason for all the difficulties. In any case, the lake was rough today and drilling could not take place until the weather calmed down.
- (inserted: 17.12.2010 22:00)
The lake was a bit wavy this morning, but has ...The lake was a bit wavy this morning, but has calmed down. Recovery in the top section of Hole C has been low due to a number of factors. After loosing the bottom-hole assembly, the drillers were left with one tool liner. A decision was made by head PI Moti Stein not to risk bending the last one on an HPC shot (which could have resulted in a delay of over a week while we waited for the spare parts to arrive), but rather go ahead with the Alien tool from the start. Moti, who is out on the rig today (as he is most days) reports that they have now penetrated a sequence of muds and white laminae, which he hopes is aragonite.
- (inserted: 15.12.2010 22:00)
Well, we moved to the second hole (Hole B) and ...Well, we moved to the second hole (Hole B) and started coring. The first HPC shot somehow bent the inner-tube of the tool so the drillers switched to the Alien. Recovery has been bad. Out of 21.3 m drilled, only 4.2 were recovered. Beau suggested that chunks of hard salt may be getting caught in the bottom of the pipe (at the drill-bit) causing it to act like a solid bit rather than a coring tool - i.e. it is just drilling away at the muds and clays, without allowing any material to enter the plastic liner. The night shift called to say that they could not latch on to their tool. After tripping up, it was discovered that a weld broke and the entire bottom-hole assembly was lost in the hole. On to Hole C...
- (inserted: 14.12.2010 22:00)
We have begin coring again! Last night's night shift moved ...We have begin coring again! Last night's night shift moved the platform over by about 5 meters and tripped down the pipe. The day shift today started coring the second hole (2 runs with the HPC and then drilling with the Alien was the plan). The logging crew are still here and will wait to log at least part of the new hole. The weather has been good. Just a little wind and some waves this morning. It is good to get back to work.
- (inserted: 12.12.2010 22:00)
We are back on the rig! After a very bad ...We are back on the rig! After a very bad storm (the worst on the Dead Sea since 1980) we sent out the night shift last night long with the loggers. The platform was in pretty good shape except for the back end of the drill rig, which had moved (pivoted) by about 50 cm and onto 2 hydraulic lines. Waves? That's 40,000 lbs to move... The day shift today should get the logging finished and the night shift will trip up, move over a few meters and trip down to the new hole. They may get one or two cores done before the day shift, but we should be back in full swing tomorrow morning.